OpenStudy (anonymous):

y/9=4-9x....y/9=4x-9/x ... now im stuck. I need to solve for y.

OpenStudy (anonymous):

I tried Mulitplying both sides by 1/9, but I dont think that would work

OpenStudy (anonymous):

multiply both sides with 9

OpenStudy (anonymous):

and then solve further

OpenStudy (anonymous):

I got 36x-81/9x

OpenStudy (anonymous):

wait i will solve it send yu a pic

OpenStudy (anonymous):

y/9=4-9x y = 9*(4-9x) y = 36 - 81x substitute this to ..y/9=4x-9/x y/9=4x-9/x y = 9*(4x-9/x) y = 36x - 81/x y = (36x^2 - 81)/x xy = 36x^2 - 81 x(36 - 81x) = 36x^2 - 81 36x - 81x^2 = 36x^2 - 81 36x = 36x^2 +81x^2 - 81 36x = 117x^2 - 81 0 = 117x^2 -36x -81

OpenStudy (anonymous):

then

OpenStudy (anonymous):

@sheetalvee in the 2nd equation is the whole thing on top of x or only 9?

Nnesha (nnesha):

just to explain that y is already multiply by 1/9 that's why its y/9 so that's why you multiplying both side 9

OpenStudy (anonymous):

wait, okay I need to solve for y right, then after I need to differentiate to get y1

OpenStudy (anonymous):

117x^2 -36x -81 = 0 9(13x^2 -4x -9) = 0

OpenStudy (anonymous):

117(x -(2/13))^2 - (1089/13) = 0

OpenStudy (anonymous):

\[117(x - \frac{ 2 }{ 13 })^2 - \frac{ 1089 }{ 13 } = 0\]

OpenStudy (anonymous):

oh pellet, Guys Im sorry, It wa 9/y = 4x-9 / x

OpenStudy (anonymous):

Thats why I said mulitply both sides by 1/9

OpenStudy (anonymous):

So \[x = - {\frac{ 9 }{ 13 }}\] \[x = 1\]

OpenStudy (anonymous):

Now Im just confusing myself. Ill send you the actual question

OpenStudy (anonymous):

hmmm :(

OpenStudy (anonymous):

\Part b

OpenStudy (anonymous):

you can go ahead and solve it for \(y\) in maybe 3 steps

OpenStudy (anonymous):

I got up to this step: 9/y=4x-9/x

OpenStudy (anonymous):

\[\frac{9}{x}+\frac{9}{y}=4\\ \frac{9}{y}=4-\frac{9}{x}=\frac{4x-9}{x}\]

OpenStudy (anonymous):

then flip it and get \[\frac{y}{9}=\frac{x}{4x-9}\]

OpenStudy (anonymous):

multiply by 9 and get \[y=\frac{9x}{4x-9}\]

OpenStudy (anonymous):

find the derivative via the annoying but useful quotient rule

OpenStudy (anonymous):

9/y = 4 - (9/x) y = 9/(4 - (9/x)) y = 9 * (x/(4x-9)) = 9x/(4x -9)

OpenStudy (anonymous):

how is that webassign anyways?

OpenStudy (anonymous):

wasnt i suppose to multiply 4x-9 by 9 aswell ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why not

OpenStudy (anonymous):

\[\frac{x}{9}=\frac{4}{5}\\ x=\frac{9\times 4}{5}\]

OpenStudy (anonymous):

you are not building up a fraction by multiplying top and bottom by some number you are multiplying a fraction by a number

OpenStudy (anonymous):

multiply means multiply in the numerator, not multiply top and bottom

OpenStudy (anonymous):

Ohhhhhhhhhh yesss. Thank you !

OpenStudy (anonymous):

and what was the question about webassign

OpenStudy (anonymous):

\[\frac{ d }{ dx } \left( \frac{ 9x }{ 4x-9 } \right) = -{\left( \frac{ 81 }{ (9-4x)^2} \right)}\]

OpenStudy (anonymous):

the question was, how is that webassign? easy? hard? annoying??

OpenStudy (anonymous):

Um it's okay. The questions are similar to our textbook so it's not bad as long as we finish the textbook questions. But Im doing it together which I realize now I shouldn't. Overall it should be easy if I had done the textbook questions first

OpenStudy (anonymous):

its simple all you need to do is bring both the equations in standard form that is you should isolate y or x in both the equations then simply substitute the vale of any ne equation into the other and solve for y-IF U KNOW WHAT I MEAN

OpenStudy (anonymous):

its a trick question i got it

OpenStudy (anonymous):

hold on for a sec i will upload a pic

OpenStudy (anonymous):

@sheetalvee okay i got yuo

OpenStudy (anonymous):

@sheetalvee here's what i think you see both equation =y/9 so you can set both part on right side equal to each other and solve for x and then substitute your value for x in any one of the original equations to get y

OpenStudy (anonymous):

@sheetalvee let me know what you think

OpenStudy (anonymous):

Yep I got it ! :)